OpenStudy (anonymous):

Simplify the following expression; express your answer as a power with a base of 5. (244140625 x 48828125^3)^4 x 625^8 ________________________________________ (9765625^4 x 15625^3 ÷ 78125^8)^6

OpenStudy (anonymous):

Thankfully these all end in 5's or 0's so they all have 5 as a factor (probably several times over seeing how ginormous they are). You need to reduce these FIRST before trying to type that all in. Do this by taking log\(\_5\)(put # here)

OpenStudy (anonymous):

\[\huge log_5 (number )\]

OpenStudy (anonymous):

This will tell you the power for each \[\huge \log_5 (625) = 4 \] meaning \[\huge 5^4=625\]

OpenStudy (anonymous):

Phew this took a bit but here you go: \[\huge \frac{(5^{12} * (5^{11})^3)^4*(5^4)^8}{((5^{10})^4*(5^6)^3/(5^7)^8)^6} \] You should be able to find the answer from here :-) Just review your rule of exponents: http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

After simplification, you get \[ \frac{5^{(45) 4+32}}{5^{(58-56) 6}}=\frac { 5^{212}}{5^{12}}=5^{200} \]